# ProtoTS: Learning Hierarchical Prototypes for Explainable Time Series Forecasting

### v3

@@ 3.3 Hierarchical Prototype Learning
While the above describes a two-level hierarchy, the structure is general: any leaf prototype can be incrementally split to form a multi-level tree. However, not all prototypes need splitting, and some splits may yield overly similar temporal patterns. Therfore, we propose a splitting rule for determining which leaf prototypes should be further refined. The rule is designed to identify prototypes whose current temporal patterns are insufficiently representative of their associated instances and thus require additional fine-grained refinement. The pseudocode is provided in Algorithm 1.
@@ 4.1 Overall Performance
Datasets Load Forecasting dataset (LOF) is a synthetic electric load dataset, including 22 supporting covariates and covering four regions. We also evaluated on the Electricity Price Forecasting (EPF) dataset (lago2021forecasting), which covers hourly electric prices from five Nordic markets. The specific details of the datasets can be found in the Appendix B.1.
Baselines To provide a wide‑ranging comparison, we benchmark eight models in total: five state-of-the-art deep-learning methods (TimeXer (wang2024timexer), iTransformer (liu2023itransformer), TiDE (das2023long), TFT (lim2021temporal), N‑BEATSx (olivares2023neural)) together with three strong machine-learning baselines (XGBoost (chen2016xgboost), LightGBM (ke2017lightgbm), pyGAM (serven2018pygam)). All deep-learning models support exogenous variables as inputs, and we further extend TimeXer and iTransformer to incorporate future covariates. The machine learning models generate forecasts recursively with exogenous inputs.
@@ 4.2 Ablation Study and Sensitivity Analysis
Proportion of Train Data To evaluate the impact of limited training data on model performance, we gradually increase the size of the PC dataset from 50% to 100%. As the available data decreased, existing baselines such as TimeXer, iTransformer, and TiDE show clear drops in MSE. In contrast, ProtoTS maintains stable results with only minor performance loss, which improves high data efficiency. The comparison results are illustrated in Figure 3(b).
@@ 4.3 Quantitative Evaluation of Interpretability
In this section, we conduct a quantitative evaluation of interpretability. We aim to answer the two questions: 1) How understandable and accurate are the prototypes in explaining the predictions? 2) How usable are the explanation systems provided by different interpretable models (ProtoTS, TFT, NBEATSx)? Specifically, we provide 24 users with different explanations as well as the corresponding predictions, where the three models are presented randomly to avoid bias. Users were asked concrete questions about the input variables, such as "Given this prediction and explanations, which season does the predicted day most likely belong to?". To evaluate interpretability, we define User Precision (UPrec) as the proportion of correctly answered variable-related questions, reflecting whether users can accurately infer the values of relevant input features from the explanations. To further assess usability, we adopt the System Usability Scale (SUS) (brooke1996sus), a widely used questionnaire that measures users’ perceived ease of use and overall satisfaction with a system. The quantitative results are reported in Table 4, with additional details in Appendix D.
@@ 5 Conclusion
Ethics statement
We provide an ethics statement as follows:
  • Human subjects research: Our experiments include a user study (see Section 4.3). The study does not raise any ethical concerns, and all participants gave informed consent prior to participation.
  • Datasets: All datasets used in our work are publicly available. The released dataset is a simulation of real-world data and does not contain any actual sensitive information.
  • Fairness and bias: Our work does not raise fairness or bias concerns.
Reproducibility statement
We confirm that the contents of the main text and the appendix are sufficient to reproduce our work and our experiments are reproducible. Section 3 clearly describes the proposed method and workflow. Appendix B.1 provides detailed dataset descriptions, and Appendix B.2 provides sufficient implementation details (including hyperparameters and training settings). We also include in the abstract a link to an anonymous repository containing our source code and the released dataset; the repository provides scripts for data preparation, training, and evaluation, along with configuration files and random seeds. We have verified that the code is consistent with the descriptions in the paper and that the reported experiments can be reproduced following the provided instructions.
@@ B.1 Dataset Descriptions
LOF dataset The LOF dataset is an synthetic electric load forecasting dataset. It is designed to predict future electricity demand at the provincial and city levels, enable electricity generators, distributors, and suppliers to plan effectively ahead and promote energy conservation among the users (nti2020electricity). Due to the influence of human industrial and commercial activities, electricity demand exhibits a clear daily cycle, maintaining a relatively stable and predictable pattern under normal conditions. The difficulty of load forecasting mainly arises from special instances that are strongly correlated with a large amount of exogenous information, such as time-related attributes and comprehensive weather data. Consequently, accurate electricity load forecasting requires models to not only effectively handle these variables but also provide insights into how they affect prediction results. Hence, electricity load forecasting is a typical task of time series forecasting with exogenous variables, characterized by a high demand for interoperability. The LOF dataset ranges across three years, with a sampling interval of 15 minutes. The LOF dataset includes 22 exogenous variables, consisting of 14 discrete time-related variables and 8 continuous weather variables. The specific details of these exogenous variables are presented in Table 10.
EPF dataset The EPF dataset is an electricity price forecasting datasets (lago2021forecasting). It collects from five distinct day-ahead electricity markets in Northern Europe, covering a six year period with a sampling interval of 1 hour. The five datasets are described as follows: (1) NP corresponds to the Nord Pool market, providing hourly electricity prices, grid load data, and wind power forecasts from 2013-01-01 to 2018-10-24. (2) PJM refers to the Pennsylvania-New Jersey-Maryland market, including zonal electricity prices for the Commonwealth Edison (COMED) zone, along with system-wide load and COMED-specific load forecasts, covering the period from 2013-01-01 to 2018-10-24. (3) BE represents the Belgian electricity market, offering hourly price data, load forecasts for Belgium, and generation forecasts for France, collected from 2011-01-09 to 2016-10-31. (4) FR pertains to the French electricity market, comprising hourly price records, load forecasts, and generation forecasts over the timeframe from 2012-01-09 to 2017-10-31. (5) DE covers the German electricity market, documenting hourly electricity prices, zonal load forecasts for the TSO Amprion area, as well as wind and solar generation forecasts, with data available from 2012-01-09 to 2017-10-31.
@@ B.2 Implementation Details
All experiments are implemented using PyTorch (paszke2019pytorch) and executed on a single NVIDIA V100 GPU with 32GB of memory. For model optimization, we employ the Adam (kingma2014adam) optimizer with an initial learning rate of and use our Loss as the objective function. The training is set for a maximum of 30 epochs, with early stopping applied to prevent overfitting. In our proposed model, the number of bottleneck fusion blocks is selected from the set {1, 2, 3}, while the dimension of the series representations, denoted as model dimensions, is chosen from {32, 64, 128, 256, 512}, the number of initial prototypes is searched from {6, 8, 12}, the deepest level is chosen from {1, 2, 3}, the number of each prototype can split is selected from the set {2, 3}. All baseline models used for comparison are re-implemented based on the TimeXer Repository benchmark. All baseline use REVIN (kim2021reversible) in training and test process.
@@ Appendix D User Study Details
We conducted the user study using a questionnaire format, the specific layout of which is shown in Figure 5. The questionnaire consists of two parts: Variable-related Questionnaire and System Usability Scale (SUS) Questionnaire. The first part is Variable-related Questionnaire, aims to evaluate the accuracy and comprehensibility of explanations provided by different explainable models. Participants are asked to perform causal inference on variables—such as holidays, seasons, Chinese New Year, and other related variables—based on the model’s predictions and its generated explanations. This section has unique, objective correct answers, and we measure the readability of the explanations by calculating the accuracy of participants’ responses. The second part is System Usability Scale (SUS) Questionnaire, designed to assess the usability of the explanation systems offering different levels of explainability. Participants are asked to subjectively rate the system based on their experience completing the first part. The SUS yields a single number representing a composite measure of the overall usability of the system under study. Note that scores for individual items are not meaningful on their own. To calculate the SUS score, first sum the score contributions from each item. Each item’s contribution ranges from 0 to 4. For items 1, 3, 5, 7, and 9, the contribution is the scale position minus 1. For items 2, 4, 6, 8, and 10, the contribution is 5 minus the scale position. Multiply the total sum of these contributions by 2.5 to obtain the overall SUS score, which ranges from 0 to 100.
@@ Appendix F Discussion
  • Dependence on domain expertise for refinement: The interpretability of ProtoTS benefits from expert-guided prototype editing. However, this process requires sufficient domain knowledge and manual effort. In domains where expert resources are hard to access or knowledge is difficult to formalize, the usability and effectiveness of this feature may be limited.
@@ Appendix G Multivariate Time Series Forecasting
  • ETTh1 and ETTh2 (zhou2021informer): Hourly electricity load data from two different nodes in a power grid. They capture seasonal patterns and variable interactions across multiple features, serving as standard benchmarks for long-term forecasting.
  • ETTm1 and ETTm2 (zhou2021informer): Minute-level electricity load datasets focusing on finer temporal granularity. They emphasize short-term dynamics and present challenges in modeling high-frequency fluctuations.
  • Weather (zhou2021informer): A multivariate weather dataset containing temperature, humidity, wind speed, and other meteorological variables collected from multiple stations.
  • Exchange (lai2018modeling): A financial time series dataset with daily exchange rates of eight foreign currencies.
@@
Baseline We compare ProtoTS against a broad set of baselines, including: Transformer-based models (TimeXer (wang2024timexer), iTransformer (liu2023itransformer), PatchTST (nie2022time), TimesNet (wu2022timesnet), Crossformer (zhang2023crossformer), Stationary (liu2022non), Autoformer (wu2021autoformer)), MLP-based architectures (TiDE (das2023long), DLinear (zeng2023transformers), RLinear (li2023revisiting), SCINet (liu2022scinet)).
@@
Table 11: Full results of multivariate time series forecasting task, where the input window is set to 96, and the forecast window is set to {96, 192, 336, 720}. Results are citepd from TimeXer (wang2024timexer), otherwise reproduced. $1^{\mathrm{st}}$ denotes the first count.

### v4

@@ 3.3 Hierarchical Prototype Learning
While the above describes a two-level hierarchy, the structure is general: any leaf prototype can be incrementally split to form a multi-level tree. However, not all prototypes need splitting, and some splits may yield overly similar temporal patterns. Therefore, we propose a splitting rule for determining which leaf prototypes should be further refined. The rule is designed to identify prototypes whose current temporal patterns are insufficiently representative of their associated instances and thus require additional fine-grained refinement. The pseudocode is provided in Algorithm 1.
@@ 4.1 Overall Performance
Datasets Load Forecasting dataset (LOF) is a synthetic electric load dataset, including 22 supporting covariates and covering four regions. We also evaluated on the Electricity Price Forecasting (EPF) dataset (Lago et al., 2021), which covers hourly electric prices from five Nordic markets. The specific details of the datasets can be found in the Appendix B.1.
Baselines To provide a wide‑ranging comparison, we benchmark eight models in total: five state-of-the-art deep-learning methods (TimeXer (Wang et al., 2024), iTransformer (Liu et al., 2023), TiDE (Das et al., 2023), TFT (Lim et al., 2021), N‑BEATSx (Olivares et al., 2023)) together with three strong machine-learning baselines (XGBoost (Chen and Guestrin, 2016), LightGBM (Ke et al., 2017), pyGAM (Servén and Brummitt, 2018)). All deep-learning models support exogenous variables as inputs, and we further extend TimeXer and iTransformer to incorporate future covariates. The machine learning models generate forecasts recursively with exogenous inputs.
@@ 4.2 Ablation Study and Sensitivity Analysis
Proportion of Train Data To evaluate the impact of limited training data on model performance, we gradually increase the size of the PC dataset from 50% to 100%. As the available data decreased, existing baselines such as TimeXer, iTransformer, and TiDE show clear drops in MSE. In contrast, ProtoTS maintains stable results with only minor performance loss, which improves high data efficiency (Jin et al., 2023b). The comparison results are illustrated in Figure 3(b).
@@ 4.3 Quantitative Evaluation of Interpretability
In this section, we conduct a quantitative evaluation of interpretability. We aim to answer the two questions: 1) How understandable and accurate are the prototypes in explaining the predictions? 2) How usable are the explanation systems provided by different interpretable models (ProtoTS, TFT, NBEATSx)? Specifically, we provide 24 users with different explanations as well as the corresponding predictions, where the three models are presented randomly to avoid bias. Users were asked concrete questions about the input variables, such as "Given this prediction and explanations, which season does the predicted day most likely belong to?". To evaluate interpretability, we define User Precision (UPrec) as the proportion of correctly answered variable-related questions, reflecting whether users can accurately infer the values of relevant input features from the explanations (Lee et al., 2025). To further assess usability, we adopt the System Usability Scale (SUS) (Brooke and others, 1996), a widely used questionnaire that measures users’ perceived ease of use and overall satisfaction with a system. The quantitative results are reported in Table 4, with additional details in Appendix D.
@@ 5 Conclusion
Acknowledgements
This work was partly supported by the National Natural Science Foundation of China (NSFC) (NO. 62476279, NO. U2436209), Major Innovation & Planning Interdisciplinary Platform for the “Double-First Class” Initiative, Renmin University of China, the Fundamental Research Funds for the Central Universities, and the Research Funds of Renmin University of China No.24XNKJ18. This work was partially supported by fund for building world-class universities (disciplines) of Renmin University of China and Public Computing Cloud, Renmin University of China. This work was partially supported by DAMO Academy (Hupan Laboratory) through DAMO Academy (Hupan Laboratory) Innovative Research Program.
@@ B.1 Dataset Descriptions
LOF dataset The LOF dataset is an synthetic electric load forecasting dataset. It is designed to predict future electricity demand at the provincial and city levels, enable electricity generators, distributors, and suppliers to plan effectively ahead and promote energy conservation among the users (Nti et al., 2020). Due to the influence of human industrial and commercial activities, electricity demand exhibits a clear daily cycle, maintaining a relatively stable and predictable pattern under normal conditions. The difficulty of load forecasting mainly arises from special instances that are strongly correlated with a large amount of exogenous information, such as time-related attributes and comprehensive weather data. Consequently, accurate electricity load forecasting requires models to not only effectively handle these variables but also provide insights into how they affect prediction results. Hence, electricity load forecasting is a typical task of time series forecasting with exogenous variables, characterized by a high demand for interoperability. The LOF dataset ranges across three years, with a sampling interval of 15 minutes. The LOF dataset includes 22 exogenous variables, consisting of 14 discrete time-related variables and 8 continuous weather variables. The specific details of these exogenous variables are presented in Table 10.
EPF dataset The EPF dataset is an electricity price forecasting datasets (Lago et al., 2021). It collects from five distinct day-ahead electricity markets in Northern Europe, covering a six year period with a sampling interval of 1 hour. The five datasets are described as follows: (1) NP corresponds to the Nord Pool market, providing hourly electricity prices, grid load data, and wind power forecasts from 2013-01-01 to 2018-10-24. (2) PJM refers to the Pennsylvania-New Jersey-Maryland market, including zonal electricity prices for the Commonwealth Edison (COMED) zone, along with system-wide load and COMED-specific load forecasts, covering the period from 2013-01-01 to 2018-10-24. (3) BE represents the Belgian electricity market, offering hourly price data, load forecasts for Belgium, and generation forecasts for France, collected from 2011-01-09 to 2016-10-31. (4) FR pertains to the French electricity market, comprising hourly price records, load forecasts, and generation forecasts over the timeframe from 2012-01-09 to 2017-10-31. (5) DE covers the German electricity market, documenting hourly electricity prices, zonal load forecasts for the TSO Amprion area, as well as wind and solar generation forecasts, with data available from 2012-01-09 to 2017-10-31.
@@ B.2 Implementation Details
All experiments are implemented using PyTorch (Paszke, 2019) and executed on a single NVIDIA V100 GPU with 32GB of memory. For model optimization, we employ the Adam (Kingma, 2014) optimizer with an initial learning rate of and use our Loss as the objective function. The training is set for a maximum of 30 epochs, with early stopping applied to prevent overfitting. In our proposed model, the number of bottleneck fusion blocks is selected from the set {1, 2, 3}, while the dimension of the series representations, denoted as model dimensions, is chosen from {32, 64, 128, 256, 512}, the number of initial prototypes is searched from {6, 8, 12}, the deepest level is chosen from {1, 2, 3}, the number of each prototype can split is selected from the set {2, 3}. All baseline models used for comparison are re-implemented based on the TimeXer Repository benchmark. All baseline use REVIN (Kim et al., 2021) in training and test process.
@@ Appendix D User Study Details
We conducted the user study using a questionnaire format, the specific layout of which is shown in Figure 5. The questionnaire consists of two parts: Variable-related Questionnaire and System Usability Scale (SUS) Questionnaire. The first part is Variable-related Questionnaire, aims to evaluate the accuracy and comprehensibility of explanations provided by different explainable models. Participants are asked to perform causal inference on variables—such as holidays, seasons, Chinese New Year, and other related variables—based on the model’s predictions and its generated explanations. This section has unique, objective correct answers, and we measure the readability of the explanations by calculating the accuracy of participants’ responses. The second part is System Usability Scale (SUS) Questionnaire, designed to assess the usability of the explanation systems offering different levels of explainability. Participants are asked to subjectively rate the system based on their experience completing the first part (Xu et al., 2024). The SUS yields a single number representing a composite measure of the overall usability of the system under study. Note that scores for individual items are not meaningful on their own. To calculate the SUS score, first sum the score contributions from each item. Each item’s contribution ranges from 0 to 4. For items 1, 3, 5, 7, and 9, the contribution is the scale position minus 1. For items 2, 4, 6, 8, and 10, the contribution is 5 minus the scale position. Multiply the total sum of these contributions by 2.5 to obtain the overall SUS score, which ranges from 0 to 100.
@@ Appendix F Discussion
  • Dependence on domain expertise for refinement: The interpretability of ProtoTS benefits from expert-guided prototype editing. However, this process requires sufficient domain knowledge and manual effort. In domains where expert resources are hard to access or knowledge is difficult to formalize, the usability and effectiveness of this feature may be limited (Zhang et al., 2024a).
@@ Appendix G Multivariate Time Series Forecasting
  • ETTh1 and ETTh2 (Zhou et al., 2021): Hourly electricity load data from two different nodes in a power grid. They capture seasonal patterns and variable interactions across multiple features, serving as standard benchmarks for long-term forecasting.
  • ETTm1 and ETTm2 (Zhou et al., 2021): Minute-level electricity load datasets focusing on finer temporal granularity. They emphasize short-term dynamics and present challenges in modeling high-frequency fluctuations.
  • Weather (Zhou et al., 2021): A multivariate weather dataset containing temperature, humidity, wind speed, and other meteorological variables collected from multiple stations.
  • Exchange (Lai et al., 2018): A financial time series dataset with daily exchange rates of eight foreign currencies.
@@
Baseline We compare ProtoTS against a broad set of baselines, including: Transformer-based models (TimeXer (Wang et al., 2024), iTransformer (Liu et al., 2023), PatchTST (Nie et al., 2022), TimesNet (Wu et al., 2022), Crossformer (Zhang and Yan, 2023), Stationary (Liu et al., 2022b), Autoformer (Wu et al., 2021)), MLP-based architectures (TiDE (Das et al., 2023), DLinear (Zeng et al., 2023), RLinear (Li et al., 2023b), SCINet (Liu et al., 2022a)).
@@
Table 11: Full results of multivariate time series forecasting task, where the input window is set to 96, and the forecast window is set to {96, 192, 336, 720}. Results are citepd from TimeXer (Wang et al., 2024), otherwise reproduced. $1^{\mathrm{st}}$ denotes the first count.